40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 8668     Entry time: Tue Jun 4 10:33:09 2013
Author: rana 
Type: Update 
Category: SUS 
Subject: ETMY oplev path redone 

 We also redid the oplev path.

Someone had used the illegal tactic of using aluminum dogs to hold down mounts which already have screw holes in the bases. This is too flimsy to use.

We then adjusted the position of the second lens to get a ~1 mm spot on the QPD. There is a lot of stray red light from extra reflections, but so be it.

Its possible that we are now clipping the IPANG path, but we will need to lock the Y arm and verify the input pointing to be sure. At this point we are ready to measure the ETMY OL loop gain and tune, etc.

ELOG V3.1.3-